OpenStudy (anonymous):

if tan@=a/b find sin2+cos2

myininaya (myininaya):

sin(2) is irrational cos(2) is also irrational

myininaya (myininaya):

the sum of sin(2) and cos(2) is also irrational this is something you plug into the calculator and approximate

myininaya (myininaya):

we don't need the if part

myininaya (myininaya):

unless you typed something wrong

OpenStudy (anonymous):

maybe it was tan(2)? since theres a @ there.

OpenStudy (anonymous):

this is not 2 that is square

OpenStudy (anonymous):

so you want:\[\sin^2(\theta)+\cos^2(\theta)\]?

OpenStudy (anonymous):

yes you are right

myininaya (myininaya):

this is identity sin^2(theta)+cos^2(theta)=1

myininaya (myininaya):

we don't need the if part lol

OpenStudy (anonymous):

that is sin square theta plus cos square theta

OpenStudy (anonymous):

right. That is a trig identity, and its equal to 1, just like Myininaya said.

OpenStudy (anonymous):

ok sir how to solve that problem?

myininaya (myininaya):

?

OpenStudy (anonymous):

if tan @ = a/b find sin square theta + cos square theta

OpenStudy (anonymous):

its not really something to solve, its a fact. something you should have memorized. the question could have been: If the sky is black, and cheese grows old after 5 days, what is:\[\sin^2(\theta)+\cos^2(\theta)\] and the answer would still be 1. Its because of unit circle.

myininaya (myininaya):

do you know the Pythagorean thm?

OpenStudy (anonymous):

yes

myininaya (myininaya):